OpenStudy (anonymous):

If the sum to infinity of a G.P. is 81/4 and its second term is -9, the common ratio is....?

OpenStudy (shubhamsrg):

Sum of infinite terms of GP is finite only if its common ratio is a fraction i.e. |r| < 1 You must be familiar with the sum of infinite terms of a GP with fraction as common ratio ?

OpenStudy (anonymous):

this's my formula \[\frac{ -9\div r }{ 1-r }=\frac{ 81 }{ 4 }\]

OpenStudy (shubhamsrg):

that is right

OpenStudy (anonymous):

really? but i cannot solve it

OpenStudy (shubhamsrg):

You'll get a quadratic equation in r . Where are you stuck?

OpenStudy (anonymous):